OpenStudy (anonymous):

Kimberly has a bag of candy consisting of 9 red sweets, 15 orange and 8 green. If he eats 2 of the sweets one after the other, what is the probability that the following will occur: a) the first sweet eaten was orange and the second red b) both sweets were green (Please show work so I can see how you got the answer) Thanks in advance!

OpenStudy (anonymous):

Eating one will change the number in the bag and the number of that color, as this is sampling without replacement. Total N= 32, then 31. p(orange, then red) = p(orange) p(red) = (15/32)(9/31) p(green, green) = (8/32)(7/31)

OpenStudy (anonymous):

Thank you so much!